# A Theory of Semantic Communication

### v3

@@ Appendix B Properties of the Six Subsets
$\underline{D}(L^{\prime})+{\underline{J}}^{({\underline{t}}^{\prime})}\left(L_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-L^{\prime}\right)$
$+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}{{\underline{J}}^{\left({\underline{t}}^{\prime}\right)}\left(L_{{\underline{{\bm{U}}}}^{(t)}}\!-\!L_{{\underline{{\bm{U}}}}^{\left(t-1\right)}}\right)}+{\underline{J}}^{\left({\underline{t}}^{\prime}\right)}\left(L^{\prime\prime}\!-\!L_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}\!-\!1\right)}}\right)$
$\leq\tilde{D}(L^{\prime\prime})$
$\leq\underline{D}(L^{\prime})+{\underline{J}}^{(\underline{t}^{\prime\prime})}\left(L_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-L^{\prime}\right)$
$+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}{{\underline{J}}^{\left(\underline{t}^{\prime\prime}\right)}\left(L_{{\underline{{\bm{U}}}}^{(t)}}\!-\!L_{{\underline{{\bm{U}}}}^{\left(t-1\right)}}\right)}+{\underline{J}}^{\left(\underline{t}^{\prime\prime}\right)}\!\left(L^{\prime\prime}\!-\!L_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}\!-\!1\right)}}\right).$
@@
Proof. We first prove that $\underline{D}(L)$ is convex. It is sufficient to prove that for any $L^{\prime}$, $L^{\prime\prime}\in[L_{\text{min}},L_{\text{max}}]$ and $\lambda\in(0,1)$, we have
@@
$\underline{D}\left(\hat{L}\right)\leq\hat{D}\leq\overline{D}\left(\hat{L}\right).$
@@
$\underline{D}(L^{\prime})\leq D^{\prime}\text{ and }\underline{D}(L^{\prime\prime})\leq D^{\prime\prime}.$
@@
$\begin{split}\underline{D}\left(\hat{L}\right)&\leq\lambda\underline{D}(L^{\prime})+\left(1-\lambda\right)\underline{D}(L^{\prime\prime})\\ &\leq\lambda D^{\prime}+\left(1-\lambda\right)D^{\prime\prime}=\hat{D}.\end{split}$
Likewise, we can also prove that $\hat{D}\leq\overline{D}\left(\hat{L}\right)$. The lemma is proved. $\blacksquare$
@@
Appendix D Proof of Theorem 5.6
Under the Hamming distortion measure, $\psi_{q}(\hat{w},\hat{s})$ can be refined as
$\psi_{q}(\hat{w},\hat{s})$ $=$ $\sum_{w,s}q(w)p(s|w)c(\hat{s}|s)d(w,\hat{w})$
$=$ $\sum_{w,s}q(w)p(s|w)c(\hat{s}|s)-q(\hat{w})\sum_{s}p(s|\hat{w})c(\hat{s}|s)$
$=$ $q(\hat{s})-q(\hat{w})p(\hat{s}|\hat{w}),$
@@
$D_{\bm{P},\bm{V}^{*}_{q}}=\sum_{\hat{s}}\psi_{p}\big{(}\hat{w}_{q}(\hat{s}),\hat{s}\big{)}=1-\sum_{\hat{s}}p\big{(}\hat{w}_{q}(\hat{s})\big{)}p\big{(}\hat{s}|\hat{w}_{q}(\hat{s})\big{)}.$ (59)
An intuitive explanation of (58) and (59) is as follows. Under the prior distribution $q(w)$, (58) is simply the maximum a posteriori (MAP) decoding for a given $\hat{s}$. When the decoding scheme is determined, a received message $\hat{s}$ is decoded to $\hat{w}_{q}(\hat{s})$. As a result, the probability of successful decoding is the summation $\sum_{\hat{s}}p\big{(}\hat{w}_{q}(\hat{s})\big{)}p\big{(}\hat{s}|\hat{w}_{q}(\hat{s})\big{)}$ with each term being the probability of transmitting $\hat{w}_{q}(\hat{s})$ and receiving $\hat{s}$. The Hamming distortion, i.e., the semantic error, is then given by (59).
@@
$D_{\bm{P},\bm{V}^{*}_{p}}=1-\sum_{\hat{s}}p\big{(}\hat{w}_{p}(\hat{s})\big{)}p\big{(}\hat{s}|\hat{w}_{p}(\hat{s})\big{)}.$ (61)
Comparing (61) and (59), $D_{\bm{P},\bm{V}^{*}_{q}}=D_{\bm{P},\bm{V}^{*}_{p}}$ if and only if (30) holds. Note that $\hat{w}_{q}(\hat{s})$ and $\hat{w}_{p}(\hat{s})$ in (60) and (58) are the solutions to the left- and right-hand sides of (30), respectively, but they are not necessarily the only solutions. Therefore, (30) does not imply $\hat{w}_{q}(\hat{s})=\hat{w}_{p}(\hat{s})$.
@@
$[s|w]=\operatorname*{arg\,min}_{s}\varphi(w,s),$
$[\hat{w}|s]=\operatorname*{arg\,min}_{w^{\prime}}\psi_{q}(w^{\prime},s)=\operatorname*{arg\,min}_{w^{\prime}}\varphi(w^{\prime},s).$
@@
$[\mathcal{S}|\mathcal{W}]=\left\{[s|w]:w\in\mathcal{W}\right\}.$
From conditions 3) and 4), we first prove that the mappings between $\mathcal{W}$ and $[\mathcal{S}|\mathcal{W}]$ and that between $[\mathcal{S}|\mathcal{W}]$ and $\mathcal{W}$ are one-to-one.
We prove this by contradiction. Suppose the mappings between $\mathcal{W}$ and $[\mathcal{S}|\mathcal{W}]$ are not one-to-one, condition 3) gives us $\left|[\mathcal{S}|\mathcal{W}]\right|>N$. On the other hand, from 4), we have $N=\left|\mathcal{W}\right|\geq[\mathcal{S}|\mathcal{W}]$. Contradiction. The mappings between $\mathcal{W}$ and $[\mathcal{S}|\mathcal{W}]$ are thus one-to-one. Likewise, it can be proven that the mappings between $[\mathcal{S}|\mathcal{W}]$ and $\mathcal{W}$ are also one-to-one.
@@
$[s|w_{(1)}]=\operatorname*{arg\,min}_{s}\varphi(w,s),$
and $\bm{V}^{*}_{q}$ maps $[s|w_{(1)}]$ to
@@
$w_{(1)}=\left[\hat{w}\left|\left[s|w_{(1)}\right]\right.\right].$
For any $1<n<N$, we assume $w_{(n)}=\left[\hat{w}\left|\left[s|w_{(n)}\right]\right.\right]$ and prove $w_{(n+1)}=\left[\hat{w}\left|\left[s|w_{(n+1)}\right]\right.\right]$ in the following.
@@
$\min_{s}\varphi\left(w_{(n+1)},s\right)=\min_{s}\varphi\left(w_{(\beta)},s\right)=$
$\varphi\left(w_{(\beta)},\left[s|w_{(n+1)}\right]\right)=\varphi\left(w_{(n+1)},\left[s|w_{(n+1)}\right]\right).$
@@
Therefore, $\beta=n+1$ and $w_{(n+1)}=\left[\hat{w}\left|\left[s|w_{(n+1)}\right]\right.\right]$.
@@
$D_{\bm{U},\bm{V}^{*}_{q}}(L)=\sum_{w,s,\hat{w}}p(w)u(s|w)v(\hat{w}|s)d(w,\hat{w})$
$=\sum_{w}p(w)d\big{(}w,\left[\hat{w}\left|\left[s|w\right]\right.\right]\big{)}=\sum_{w}p(w)\min_{w^{\prime}}d(w,w^{\prime})$
$\leq\min_{\bm{U}^{\prime},\bm{V}^{\prime}}D_{\bm{U}^{\prime},\bm{V}^{\prime}}(L),$

### v4

@@ Appendix B Properties of the Six Subsets
$\hskip-18.49411pt\underline{D}(L^{\prime})+{\underline{J}}^{({\underline{t}}^{\prime})}\left(L_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-L^{\prime}\right)$
$\hskip-18.49411pt+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}{{\underline{J}}^{\left({\underline{t}}^{\prime}\right)}\left(L_{{\underline{{\bm{U}}}}^{(t)}}\!-\!L_{{\underline{{\bm{U}}}}^{\left(t-1\right)}}\right)}+{\underline{J}}^{\left({\underline{t}}^{\prime}\right)}\left(L^{\prime\prime}\!-\!L_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}\!-\!1\right)}}\right)$
$\hskip-18.49411pt\leq\tilde{D}(L^{\prime\prime})$
$\hskip-18.49411pt\leq\underline{D}(L^{\prime})+{\underline{J}}^{(\underline{t}^{\prime\prime})}\left(L_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-L^{\prime}\right)$
$\hskip-18.49411pt+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}{{\underline{J}}^{\left(\underline{t}^{\prime\prime}\right)}\left(L_{{\underline{{\bm{U}}}}^{(t)}}\!-\!L_{{\underline{{\bm{U}}}}^{\left(t-1\right)}}\right)}+{\underline{J}}^{\left(\underline{t}^{\prime\prime}\right)}\!\left(L^{\prime\prime}\!-\!L_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}\!-\!1\right)}}\right).$
@@
Proof. We first prove that $\underline{D}(L)$ is convex. It is sufficient to prove that for any $^{\prime}, L′′[L$$^{\prime\prime}\in[L_{\text{min}},L_{\text{max}}$] and $\lambda\in(0,1)$, we have
@@
[…]
@@
[…]
@@
[…]
@@
Appendix D Proof of Proposition 5.3
[…]
$=$ $\hskip-5.69046pt\sum_{w,s}q(w)p(s|w)c(\hat{s}|s)-q(\hat{w})\sum_{s}p(s|\hat{w})c(\hat{s}|s)$
[…]
@@
$).D_{\bm{P},\bm{V}^{*}_{q}}=\sum_{\hat{s}}\psi_{p}\big{(}\hat{w}_{q}(\hat{s}),\hat{s}\big{)}=1-\sum_{\hat{s}}p\big{(}\hat{w}_{q}(\hat{s})\big{)}p\big{(}\hat{s}|\hat{w}_{q}(\hat{s}$$\big{)}$. (59)
[…]
@@
$).D_{\bm{P},\bm{V}^{*}_{p}}=1-\sum_{\hat{s}}p\big{(}\hat{w}_{p}(\hat{s})\big{)}p\big{(}\hat{s}|\hat{w}_{p}(\hat{s}$$\big{)}$. (61)
[…]
@@
$s|w]=\operatorname*{arg\,min}_{s}\varphi(w,s),$
$\hat{w}|s]=\operatorname*{arg\,min}_{w^{\prime}}\psi_{q}(w^{\prime},s)=\operatorname*{arg\,min}_{w^{\prime}}\varphi(w^{\prime},s).$
@@
$\mathcal{S}|\mathcal{W}]=\left\{[s|w]:w\in\mathcal{W}\right\}.$
From conditions 3) and 4), we first prove that the mappings between $\mathcal{W}$ and $\mathcal{S}|\mathcal{W}]$ and that between $\mathcal{S}|\mathcal{W}]$ and $\mathcal{W}$ are one-to-one.
We prove this by contradiction. Suppose the mappings between $\mathcal{W}$ and $\mathcal{S}|\mathcal{W}]$ are not one-to-one, condition 3) gives us $\left|[\mathcal{S}|\mathcal{W}]\right|>N$. On the other hand, from 4), we have $N=\left|\mathcal{W}\right|\geq[\mathcal{S}|\mathcal{W}]$. Contradiction. The mappings between $\mathcal{W}$ and $\mathcal{S}|\mathcal{W}]$ are thus one-to-one. Likewise, it can be proven that the mappings between $\mathcal{S}|\mathcal{W}]$ and $\mathcal{W}$ are also one-to-one.
@@
$s|w_{(1)}]=\operatorname*{arg\,min}_{s}\varphi(w,s),$
and $\bm{V}^{*}_{q}$ maps $s|w_{(1)}]$ to
@@
[…]
@@
$\hskip-42.67912pt\min_{s}\varphi\left(w_{(n+1)},s\right)=\min_{s}\varphi\left(w_{(\beta)},s\right)=$
$\hskip-14.22636pt\varphi\left(w_{(\beta)},\left[s|w_{(n+1)}\right]\right)=\varphi\left(w_{(n+1)},\left[s|w_{(n+1)}\right]\right).$
@@
[…]
@@
[…]
$\hskip-14.22636pt=\sum_{w}p(w)d\big{(}w,\left[\hat{w}\left|\left[s|w\right]\right.\right]\big{)}=\sum_{w}p(w)\min_{w^{\prime}}d(w,w^{\prime})$
$\hskip-14.22636pt\leq\min_{\bm{U}^{\prime},\bm{V}^{\prime}}D_{\bm{U}^{\prime},\bm{V}^{\prime}}(L),$